OpenStudy (anonymous):

order of operations 88+2 over (-1+5)^2

OpenStudy (anonymous):

BIDMAS Brackets Indices Division Multiplication Addition Subtraction

OpenStudy (anonymous):

Firstly do the sum inside the brackets, followed by the index. and then add the 88 and the 2 and do the division.

OpenStudy (anonymous):

5 5over 8

OpenStudy (anonymous):

\[\frac{ 88+2 }{ (-1+5)^2 } = \frac{ 90 }{ 4^2 } = \frac{ 90 }{ 16 }\]